What will you add to the following data : 21.5 , 34.6 , 44.8 , 11.3 , 17.9 to get a mean exactly equal to 30.35?


1
Expert's answer
2021-11-01T10:11:18-0400

Let ""x"" is added to the given set of numbers.

"\therefore \frac{21.5+34.6+44.8+11.3+17.9+x}{6}=30.35\\
\Rightarrow130.1+x=182.1\\
\Rightarrow x=52"